OpenStudy (anonymous):

Solve for A: B= 5/7 (a-9) please help!!!!!!!

OpenStudy (dumbcow):

basically use inverse operations to get "a" by itself and everything else to other side B = (5/7)(A-9) - 5/7 is being multiplied by (A-9) so to cancel this....divide (or multiply by reciprocal) (7/5)B = A-9 9 is being subtracted from A so to cancel this ...add (7/5)B +9 = A done !